Reports 1-1 of 1 Clear search Modify search
AdV-SIN (Design/Construction)
polini - 22:59 Tuesday 14 September 2021 (53149) Print this report
Characterization of new PSD responses on SQB1 moving SQB2

Last Friday I performed the angular and position scans of SQB1 to characterize the response of the new PSDs on SQB1 (logbook entry #53126).

Today I moved the angular and position degrees of freedom of SQB2 in order to characterize the response of the new PSDs on SQB1, in particular PSD3, since PSD1 is not looking at signals coming from SQB2.

The PSDs were aligned since the last measurements of Friday.

I turned off the driflt control of the cavity otherwise the incoming green beam was moving during the scans of SQB2. I re-aligned the cavity in the usual way: moving the IM and EM maximizing the flashes. After turning off the drift control, the peaks were or ~1.5V instead of 7V. To recover the alignement I took the last positions of the IM and EM, doing a trend of the data of this afternoon. I reached peaks of ~6/6.5 V.

Once in this situation, I was able to perform the scans of the angles and positions of SQB2.

The following scans have been performed:

  • TX: 2021-09-14-19h40m46-UTC    info polini       'Bench Set points:LC_TX_set [LC_TX=1239,ramp time=100]' sent to SQB2_LC
  • TY: 2021-09-14-19h45m53-UTC    info polini       'Bench Set points:LC_TY_set [LC_TY=200,ramp time=100]' sent to SQB2_LC
  • TZ: 2021-09-14-19h56m00-UTC    info polini       'Bench Set points:LC_TZ_set [LC_TZ=-2876,ramp time=100]' sent to SQB2_LC
  • X: 2021-09-14-20h04m08-UTC    info polini       'Bench Set points:LC_X_set [LC_X=200,ramp time=100]' sent to SQB2_LC
  • Y: 2021-09-14-20h22m29-UTC    info polini       'Bench Set points:LC_Y_set [LC_Y=150,ramp time=100]' sent to SQB2_LC
  • Z: 2021-09-14-20h28m36-UTC    info polini       'Bench Set points:LC_TZ_set [LC_Z=480,ramp time=100]' sent to SQB2_LC

At the end of the scans, I engaged the drift control of the filter cavity and the cavity is back locked at 7.1V.

Search Help
×

Warning

×